ANALYSIS OF THE 0_{0}^{0} AND 3_{1}^{0} BANDS IN THE Ã - X̃ TRANSITION OF YC_{2}

ROBERT R. BOUSQUET, KEI-ICHI C. NAMIKI, \underline{TIMOTHY C. STEIMLE}, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287; ANTHONY J. MERER, Department of Chemistry, University of British Columbia, Vancouver, British Columbia V6T 1Z1, Canada.

The 0_{0}^{0} and 3_{1}^{0} band systems of the Ã ^{2}A_{1} ← X̃ ^{2}A_{1} transition of YC_{2} were recorded in high resolution using laser-induced fluorescence on molecules produced in a molecular beam. Asymetry splittings in the X̃ ^{2}A_{1} state were measured by recording the optical spectrum in the presence of a weak static electric field ^{a}. Several pure rotational transitions were also recorded in the (0,0,0) X̃ ^{2}A_{1} vibronic state using pump/probe microwave optical double resonance spectroscopy. The three sets of parameters were combined to produce fine and hyperfine parameters for the Ã and X̃ states. Rotational constants and structural parameters were determined for both electronic states by fitting the combined data sets to an effective hamiltonian for a rigid molecule. An interpretation of the fine structure parameters will be given.

aR. Bousquet and T. C. Steimle, J. Chem. Phys. 114, 1306-1310 (2001)